# gFlora: a topology-aware method to discover functional co-response groups in soil microbial communities

### v1

@@ Abstract.
Microorganisms such as bacteria perform critical functions in the soil ecosystem: they mediate essential carbon and nitrogen cycling processes in soils. To manage the health and functions of soils, it is important to understand which soil functions are controlled by which microbial taxa—but this taxon-to-function link is difficult to discover because of the size and complexity of the soil ecosystem. A feasible solution is to discover functional links at the level of group instead of individual, using observational data of both taxa abundance and soil function. We thus aim to learn the functional co-response group: a group of taxa whose co-response effect (the representative characteristic of the group) associates well statistically with a functional variable. Different from the state-of-the-art method, we model the soil microbial community as an ecological co-occurrence network with the taxa as nodes (weighted by their abundance) and their relationships (a combination from both spatial and functional ecological aspects) as edges (weighted by the strength of the relationships). Then, we design a method called gFlora which notably uses graph convolution over this co-occurrence network to get the co-response effect of the group, such that the network topology is also considered in the discovery process. We evaluate gFlora on two real-world soil microbiome datasets (bacteria and nematodes) and compare it with the state-of-the-art method. gFlora outperforms this on all evaluation metrics, and discovers new functional evidence for taxa which were so far under-studied. We show that the graph convolution step is crucial to taxa with low abundance (thus removing the bias towards taxa with high abundance), and the discovered bacteria of different genera are distributed in the co-occurrence network but still tightly connected among themselves, demonstrating that topologically they fill different but collaborative functional roles in the ecological community.
@@ 1. Introduction
Instead, a feasible solution to the complex problem of microbial functional characterization is to discover not individual taxa, but functional co-response groups, and to do so by learning from observational data which measures both the state of the soil, and the microbial community. A functional co-response group is defined as a group of taxa that co-respond to functional variables and whose co-response effect (the representative characteristic of the group) remains stable and associates well statistically with the functional variable which reflects the metabolic result in the observational data (Nelson et al., 2016; Shan et al., 2023). These functional co-response groups provide not only a simpler and more interpretable functional model of the soil biology, but are also computationally more feasible, since they reduce the complexity of the problem.
@@
  • We further show that our hypothesis holds: co-occurrence network is useful for the discovery of functional co-response groups. The functionally important bacteria are distributed in the co-occurrence network (often in different clusters) and play different roles in the network. We thus see evidence that functional taxa would collaborate across clusters, likely on different aspects of the specific function.
@@ 2.1. Graph Convolution
To achieve this, we compute the adjacency matrix $\mathbf{A}\in\mathbb{R}^{p\times{p}}$ of the co-occurrence network inferred using the EleMi (Chen and Bucur, 2024) method, which infers this network from abundance data using multi-regression and was shown to be robust and provide a clearer cluster structure than other network-inference methods. In the co-occurrence network, the nodes are taxa and the undirected weighted edges are co-occurrence relationships (there are no self-loops).
@@
(1) $\mathbf{M}=\mathbf{H}\left(\widetilde{\mathbf{D}}^{-\frac{1}{2}}\widetilde{\mathbf{A}}\widetilde{\mathbf{D}}^{-\frac{1}{2}}\right)$
@@ 2.2. Functional Co-Response Group Discovery
(2) $\mathbf{s}=\mathbf{M}\mathbf{x}$
whose $i$th element is the group co-response effect in the $i$th sample. $\mathbf{x}$ represents the functional co-response group (learning goal) which is a Boolean vector with 1 as presence and 0 as absence in the group. Intuitively, $\mathbf{s}$ is a vector showing the total topological abundance of taxa in a given group $\mathbf{x}$, per sample.
@@
(4) $=\max\frac{\mathbf{x}^{T}\mathbf{M_{0}}^{T}\mathbf{y_{0}}}{\sqrt{\mathbf{x}^{T}\mathbf{M_{0}}^{T}\mathbf{M_{0}}\mathbf{x}}}$
@@
(5) $\frac{\mathbf{x}^{T}\mathbf{M_{0}}^{T}\mathbf{y_{0}}}{\sqrt{\mathbf{x}^{T}\mathbf{M_{0}}^{T}\mathbf{M_{0}}\mathbf{x}}}-\alpha\max(\|\mathbf{x}\|_{1}-k_{opt},0)$
@@
(6) $\frac{\mathbf{x}^{T}\mathbf{M_{0}}^{T}\mathbf{y_{0}}}{\sqrt{\mathbf{x}^{T}\mathbf{M_{0}}^{T}\mathbf{M_{0}}\mathbf{x}}}-\mu\|\mathbf{x}\|_{1}$
@@ 2.4. Datasets and Experimental Settings
The optimization of gFlora is based on the genetic algorithm, thus managing the search space is crucial to ensuring computational efficiency. We highly recommend reducing the number of taxa in the community to less than 500, otherwise the discovery of functional co-response groups may become too impractical due to the explosion of the search space size. The taxonomy of both bacteria and nematodes was originally assigned at the species level by consulting a database (van Rijssel et al., 2022a). We directly used nematode species as input to our algorithm, but we grouped bacterial taxonomic composition to the level of genera because the number of bacterial species is super large. To further constrain the search space and the computational complexity, we removed extremely “sparse” taxa (as suggested by EQO (Shan et al., 2023), those whose ratio of zero-abundance across soil samples is over 80% do not carry a lot of useful information in the statistical sense). Nevertheless, a lot of relatively rare taxa are still available to be chosen. After this, the dimensions of the abundance matrices of bacteria and nematodes were $125\times 164$ and $126\times 92$, respectively. The abundance matrices are normalized with the Cumulative-Sum Scaling (CSS,  (Paulson et al., 2013)) method (also as indicated in EQO (Shan et al., 2023)).
@@
In order to evaluate the effectiveness and generalization ability of our method, the data was divided into training and testing datasets (50%-50%, as in the related work (Shan et al., 2023)). The algorithm was trained on the training datasets to get the functional co-response groups. These groups were then used to compute the correlation $r$ between the group’s co-response effect and the functional variable on the corresponding testing datasets. We did not use classic cross-fold validation because of the limited sample size (we need enough samples to train and test the model). Instead, we repeated this process 100 times with the random seeds from 1 to 100 to mitigate the influence of randomness on the evaluation. Due to the sensitivity of correlation to the distribution of the data, stratified sampling on the functional variable (separated into 10 quantiles, sampled proportionately) was used in the sampling process to ensure the representativeness of the training sample.
@@
The genetic algorithm was implemented using the R package GA (version 3.2.3) (Scrucca, 2013) with the crossover probability set to 0.8, the mutation probability set to 0.1, and a population size of 200 (the same as in the related work (Shan et al., 2023)). The maximum number of iterations was set to 500. Early stop choices are also available using maximum fitness and the number of consecutive generations without any improvement in the best fitness value. Parallel computing is enabled by the R package doParallel (version 1.0.17). The penalty $\alpha$ in the original fitness function was set as the root of the maximum value of double-precision float in the current R environment (.$Machine\$double.xmax$). The $k_{opt}$ parameter (the best group size) was optimized in the range of 2 to 50 according to the AIC (The experiment is repeated 10 times for each group size with random seeds 1 to 10). $\mu$ for $\ell_{1}$ norm is optimized in the cross-validation process within $\{1/30,1/40,...,1/100\}$ on the training dataset.
To fully utilize all the data, when computing the importance of each taxon in the functional co-response group, we ran the algorithms on the entire dataset and repeated it 10 times with random seeds from 1 to 10. Besides, All statistical analyses were performed using paired two-sample t-test in R (version 4.3.1). A significance threshold of 0.05 was applied to determine statistical significance.
@@ 4. Conclusion
We show new practical insights on some under-studied taxa related to N-cycling. This is new, non-traditional knowledge, at the level of groups of microorganisms instead of individuals, which is realistic given that taxa function as part of a community. The main limitation is that our functional co-response groups are discovered based on correlations in the data, so causality remains to be validated by experimentation. However, even correlational evidence is important: since validated functional information per taxon is rarely available in the literature and the soil ecological community is complex, this evidence can serve as focused functional hypotheses to test experimentally. We also show new insight on the topological location of functionally important bacteria: they are distributed in the co-occurrence network, hinting at them collaborating across clusters of taxa.
gFlora can be easily extended to other soil functions. In the future, we anticipate developing a comprehensive soil ecological model that incorporates as many different soil functions and characteristics as possible. This model would provide a more holistic understanding of soil health, helping to balance the benefits and drawbacks of various soil management practices. Besides, both gFlora and EQO only show the linear relationship between microbiota and the functional variable. Non-linear options can also be considered in the future.

### v2

@@ Abstract.
Microorganisms such as bacteria perform critical functions in the soil ecosystem: they mediate essential carbon and nitrogen cycling processes in soils. To manage the health and functions of soils, it is important to understand which soil functions are controlled by which microbial taxa—but this taxon-to-function link is difficult to discover because of the size and complexity of the soil ecosystem. A feasible solution is to discover functional links at the level of group instead of individual, using observational data of both taxa abundance and soil function. We thus aim to learn the functional co-response group: a group of taxa whose co-response effect (the representative characteristic of the group showing the total topological abundance of taxa) co-responds (associates well statistically) to a functional variable. Different from the state-of-the-art method, we model the soil microbial community as an ecological co-occurrence network with the taxa as nodes (weighted by their abundance) and their relationships (a combination from both spatial and functional ecological aspects) as edges (weighted by the strength of the relationships). Then, we design a method called gFlora which notably uses graph convolution over this co-occurrence network to get the co-response effect of the group, such that the network topology is also considered in the discovery process. We evaluate gFlora on two real-world soil microbiome datasets (bacteria and nematodes) and compare it with the state-of-the-art method. gFlora outperforms this on all evaluation metrics, and discovers new functional evidence for taxa which were so far under-studied. We show that the graph convolution step is crucial to taxa with relatively low abundance (thus removing the bias towards taxa with higher abundance), and the discovered bacteria of different genera are distributed in the co-occurrence network but still tightly connected among themselves, demonstrating that topologically they fill different but collaborative functional roles in the ecological community.
@@ 1. Introduction
Instead, a feasible solution to the complex problem of microbial functional characterization is to discover not individual taxa, but functional co-response groups, and to do so by learning from observational data which measures both the state of the soil, and the microbial community. A functional co-response group is defined as a group of taxa that co-respond to functional variables and whose co-response effect (the representative characteristic of the group showing the total topological abundance of taxa) remains stable and associates well statistically with the functional variable which reflects the metabolic result in the observational data (Nelson et al., 2016; Shan et al., 2023). These functional co-response groups provide not only a simpler and more interpretable functional model of the soil biology, but are also computationally more feasible, since they reduce the complexity of the problem.
@@
  • We further show that our hypothesis holds: the co-occurrence network is useful for the discovery of functional co-response groups. The functionally important bacteria are distributed in the co-occurrence network (often in different clusters) and play different roles in the network. We thus see evidence that functional taxa would collaborate across clusters, likely on different aspects of the specific function.
@@ 2.1. Graph Convolution
To achieve this, we compute the adjacency matrix $\mathbf{A}\in\mathbb{R}^{p\times{p}}$ of the co-occurrence network inferred using the most recent method EleMi (Chen and Bucur, 2024). EleMi infers this network from abundance data using multi-regression. Compared with earlier methods like SPIEC-EASI, it was shown to be more robust and provide a clearer cluster structure. In the co-occurrence network, the nodes are taxa and the undirected weighted edges are co-occurrence relationships (there are no self-loops).
@@
(1) $\mathbf{M}=\mathbf{H}\left(\widetilde{\mathbf{D}}^{-\frac{1}{2}}\widetilde{\mathbf{A}}\widetilde{\mathbf{D}}^{-\frac{1}{2}}\right),$
@@ 2.2. Functional Co-Response Group Discovery
(2) $\mathbf{s}=\mathbf{M}\mathbf{x}.$
The $i$th element of $\mathbf{s}$ is the group co-response effect in the $i$th sample. $\mathbf{x}$ models the functional co-response group (which is unknown, and the learning goal): a Boolean vector of size $p$ (the number of taxa). For each taxon in $\mathbf{x}$, a value of 1 represents presence, and 0 represents absence in the group. Intuitively, $\mathbf{s}$ is a vector of size $n$ (the number of samples) showing the total topological abundance of taxa in a given group $\mathbf{x}$, per sample. With a proposed $\mathbf{x}$, the co-response effect vector $\mathbf{s}$ is computed out of this $\mathbf{x}$ and $\mathbf{M}$, which was calculated out of the taxa abundance by Eq. (1).
@@
(4) $=\max\frac{\mathbf{x}^{T}\mathbf{M_{0}}^{T}\mathbf{y_{0}}}{\sqrt{\mathbf{x}^{T}\mathbf{M_{0}}^{T}\mathbf{M_{0}}\mathbf{x}}}.$
@@
(5) $\frac{\mathbf{x}^{T}\mathbf{M_{0}}^{T}\mathbf{y_{0}}}{\sqrt{\mathbf{x}^{T}\mathbf{M_{0}}^{T}\mathbf{M_{0}}\mathbf{x}}}-\alpha\max(\|\mathbf{x}\|_{1}-k_{opt},0),$
@@
(6) $\frac{\mathbf{x}^{T}\mathbf{M_{0}}^{T}\mathbf{y_{0}}}{\sqrt{\mathbf{x}^{T}\mathbf{M_{0}}^{T}\mathbf{M_{0}}\mathbf{x}}}-\mu\|\mathbf{x}\|_{1},$
@@ 2.4. Datasets and Experimental Settings
The optimization of gFlora is based on a genetic algorithm, thus managing the search space is crucial to ensure computational efficiency. We highly recommend reducing the number of taxa in the community to less than 500, otherwise the discovery of functional co-response groups may become too impractical due to the explosion of the search space size. The taxonomy of both bacteria and nematodes was originally assigned at the species level by consulting a database (van Rijssel et al., 2022a). We directly used nematode species as input to our algorithm, but we grouped bacterial taxonomic composition to the level of genera because the number of bacterial species is very large. To further constrain the search space and the computational complexity, we removed extremely “sparse” taxa (as suggested by EQO (Shan et al., 2023), those whose ratio of zero-abundance across soil samples is over 80% do not carry a lot of useful information in the statistical sense). Nevertheless, a lot of relatively rare taxa are still available to be chosen. After this, the dimensions of the abundance matrices of bacteria and nematodes were $125\times 164$ and $126\times 92$, respectively. The abundance matrices are normalized with the Cumulative-Sum Scaling (CSS,  (Paulson et al., 2013)) method (also as indicated in EQO (Shan et al., 2023)).
@@
In order to evaluate the effectiveness and generalization ability of our method, the data was divided into training and testing datasets (50%-50%, as in the related work (Shan et al., 2023)). The algorithm was trained on the training datasets to get the functional co-response groups. These groups were then used to compute the correlation $r$ between the group’s co-response effect and the functional variable on the corresponding testing datasets. We did not use classic cross-fold validation because of the limited sample size (we need enough samples to train and test the model). Instead, we repeated this process 100 times to mitigate the influence of randomness on the evaluation. Due to the sensitivity of correlation to the distribution of the data, stratified sampling on the functional variable (separated into 10 quantiles, sampled proportionately) was used in the sampling process to ensure the representativeness of the training sample.
@@
The genetic algorithm was implemented using the R package GA (version 3.2.3) (Scrucca, 2013) with the crossover probability set to 0.8, the mutation probability set to 0.1, and a population size of 200 (the same as in the related work (Shan et al., 2023)). The maximum number of iterations was set to 500. Early stop choices are also available using maximum fitness and the number of consecutive generations without any improvement in the best fitness value. Parallel computing is enabled by the R package doParallel (version 1.0.17). The penalty $\alpha$ in the original fitness function was set as the root of the maximum value of double-precision float in the current R environment (.$Machine\$double.xmax$). The $k_{opt}$ parameter (the best group size) was optimized in the range of 2 to 50 according to the AIC (The experiment is repeated 10 times for each group size). $\mu$ for $\ell_{1}$ norm is optimized in the cross-validation process within $\{1/30,1/40,...,1/100\}$ on the training dataset.
To fully utilize all the data, when computing the importance of each taxon in the functional co-response group, we ran the algorithms on the entire dataset and repeated it 10 times. Besides, all statistical analyses were performed using paired two-sample t-test in R (version 4.3.1) with normality test. A significance threshold of 0.05 was applied to determine statistical significance.
@@ 4. Conclusion
We show new practical insights on some under-studied taxa related to N-cycling. This is new, non-traditional knowledge, at the level of groups of microorganisms instead of individuals, which is realistic given that taxa function as part of a community. The main limitation is that our functional co-response groups are discovered based on correlations in the data, so causality remains to be validated by experimentation. However, even correlational evidence is important: since validated functional information per taxon is rarely available in the literature and the soil ecological community is complex, this evidence can serve as focused functional hypotheses to test experimentally. We also show new insight on the topological location of functionally important bacteria: they are distributed in the co-occurrence network, hinting at them collaborating across clusters of taxa. However, the $Q$ value of clusters is not very high and the detection of clusters does not mean their formation is statistically separated. Therefore, this is merely an exploratory work to raise hypotheses, which need to be further validated in practice.
gFlora can be easily applied to any other measured soil functions to discover functional groups. In the future, we anticipate developing a comprehensive soil ecological model that incorporates as many soil functions as possible. With this model, we could further investigate the causal relationships between soil functions and those discovered functional groups. This model would provide a more holistic understanding of soil health, helping to balance the benefits and drawbacks of various soil management practices. Besides, both gFlora and EQO only discover the linear relationship between microbiota and the functional variable; non-linear options can also be considered in the future. Furthermore, though measures have been taken against overfitting, the sample size is relatively small thus the risk of overfitting still exists. In future studies, we would apply our model to larger datasets.